### v1

# RC-positivity, Schwarz’s lemma and comparison theorems

Abstract.
It is well-known that the classical Schwarz lemma yields an explicit comparison of two Hermitian metrics with uniform constant negative curvature bounds through holomorphic maps between complex manifolds. In this paper, we establish Schwarz lemmas for holomorphic bundle maps between abstract Hermitian holomorphic vector bundles with various positive curvature bounds. As applications, we prove Schwarz lemmas for holomorphic maps between complex manifolds whose curvature tensors are described by the notion “RC-positivity”. In particular, new diameter and volume comparison theorems are obtained by using Schwarz lemmas.
1. Introduction
The classical Schwarz lemma states that if $f:{\mathbb{D}}\rightarrow{\mathbb{D}}$ is a holomorphic map between two unit disks of the complex plane, then
(1.1) $f^{*}\omega_{\mathrm{P}}\leq\omega_{\mathrm{P}}$
where $\omega_{\mathrm{P}}$ is the Poincaré metric on ${\mathbb{D}}$. In [Ahl38], L. Ahlfors extended this result to holomorphic maps between Riemann surfaces whose curvature tensors were used in a very explicit way. The study of Schwarz’s lemma for holomorphic maps between higher dimensional complex manifolds was initiated in [Che68] by S.-S. Chern. Schwarz’s lemma was further extended by S. Kobayashi, P. Griffiths, H.-H. Wu and others. One of the most general differential geometric interpretations of Schwarz’s lemma is the following theorem established by S.-T. Yau ([Yau78]) and now it is called Yau’s Schwarz lemma. The curvature tensors involved are the Ricci curvature on the domain manifold and the holomorphic bisectional curvature (HBSC) on the target.
Theorem 1.1 (Yau).
Let $(M,\omega_{g})$ and $(N,\omega_{h})$ be two Kähler manifolds such that
(1.2) $\mathrm{Ric}(M,g)\geq-\lambda\omega_{g}\quad\mbox{and}\quad\mathrm{HBSC}(N,h)\leq-\kappa$
where $\lambda>0$ and $\kappa>0$ are two constants. If $M$ is complete and $f:M\to N$ is a non-constant holomorphic map, then
(1.3) $\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\leq\frac{\lambda}{\kappa}.$
As an application of Theorem 1.1, one deduces that there is no bounded holomorphic functions on complete Kähler manifolds with non-negative Ricci curvature, which generalizes the classical Liouville theorem on the complex plane. Moreover, there is no non-trivial holomorphic map from complete Kähler manifolds with non-negative Ricci curvature to Kähler manifolds with negative holomorphic bisectional curvature. There are many important generalizations of Yau’s Schwarz lemma. Notably, Chen-Cheng-Lu and Royden extended Yau’s result to the cases that target manifolds have negative holomorphic sectional curvature (HSC), and obtained the following result independently ([CCL79] and [Roy80]):
Theorem 1.2.
Let $(M,\omega_{g})$ and $(N,\omega_{h})$ be two Kähler manifolds such that
(1.4) $\mathrm{Ric}(M,g)\geq-\lambda\omega_{g}\quad\mbox{and}\quad\mathrm{HSC}(N,h)\leq-\kappa,$
where $\lambda>0$ and $\kappa>0$ are two constants. If $M$ is complete and $f:M\to N$ is a non-constant holomorphic map with maximal complex rank  $r_{f}$, then
(1.5) $\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\leq\frac{2r_{f}}{r_{f}+1}\frac{\lambda}{\kappa}.$
Recall that the holomorphic sectional curvature operator of $(N,h)$ is given by
(1.6) $\mathrm{HSC}(h,v)=\frac{R^{h}(v,\overline{v},v,\overline{v})}{|v|^{4}_{h}},\ \ \ v\in T^{1,0}N\setminus\{0\}.$
The holomorphic sectional curvature of $(N,h)$ has negative upper bound $-\kappa$ means
(1.7) $R^{h}(v,\overline{v},v,\overline{v})\leq-\kappa|v|^{4}_{h}.$
For further extensions of Schwarz’s lemma between complex manifolds, we refer to [MY83], [Kob98], [Zhe00], [Ni19], [YZ19], [Ni21], [Bro22] and the references therein. By adapting techniques from almost Hermitian geometry, V. Tosatti extended these results to almost Hermitian manifolds ([Tos07]), and more developments were obtained in [Kob01], [Mas21], [Yu22], [CN22] and etc.. These Schwarz lemmas have proven to be very useful in both differential geometry and algebraic geometry. For more details along this comprehensive topic, we refer to [Liu96], [LSY04], [LTXZ16], [WY16], [TY17], [DT19], [Yang21], [CLT22], [Yang24] and the references therein.
Schwarz’s lemma yields a version of comparison theorems. It is well-known that comparison theorems are important tools for understanding geometric concepts in differential geometry. For complete Riemannian manifolds $(M,g)$ with Ricci curvature $\mathrm{Ric}(g)\geq(n-1)g$, Myers [Mye41] established the diameter comparison theorem and Cheng [Che75] obtained the diameter rigidity theorem. The Bishop-Gromov volume comparison theorem (e.g. [BC64], [Gro07], [CE08]) asserts that $\mathrm{Vol}(M,g)\leq\mathrm{Vol}({\mathbb{S}}^{n},g_{\mathrm{can}})$, and the identity holds if and only if $(M,g)$ is isometric to the round sphere. For more details along this comprehensive topic, we refer to [CC97], [Zhu97] and [Wei07] and the references therein.
There are many notable generalied comparison theorems on Kähler manifolds. For instance, Li and Wang [LW05] obtained diameter comparison and volume comparison theorems for compact Kähler manifolds with $\mathrm{HBSC}\geq 1$, and Datar and Seshadri [DS23] established the diameter rigidity theorem. This is achieved by using Siu-Yau’s solution to the Frankel conjecture [SY80] and an interesting monotonicity formula for Lelong numbers on ${\mathbb{C}}{\mathbb{P}}^{n}$ ([Lot21]). On the other hand, by using entirely different techniques from algebraic geometry, Zhang [Zha22] obtained volume comparison and rigidity theorems under the assumption $\mathrm{Ric}(\omega)\geq(n+1)\omega$. Some interesting results are also established in [TY12], [LY18], [NZ18] and [XY24+].
All the aforementioned Schwarz lemmas are established for complex manifolds with negative curvature bounds, and comparison theorems are proved for manifolds with positive curvature bounds. In this paper, we establish Schwarz lemmas for holomorphic bundle maps between abstract Hermitian vector bundles whose curvature tensors have positive bounds described by RC-positivity. The notion RC-positivity was firstly introduced in [Yang18] by the second named author, and we refer to [Yang17], [Yang18], [Yang20] and [Yang24] for more geometric interpretations on it. As applications of these new Schwarz lemmas, we obtain several diameter and volume comparison theorems. The first main result of this paper is the following Schwarz lemma by using very weak curvature conditions.
Theorem 1.3.
Let $M$ be a compact complex manifold and $g,h$ be two Hermitian metrics. If for every $x\in M$ and every nonzero $v\in T_{x}^{1,0}M$, there exists some $w\in T_{x}^{1,0}M$ satisfying
(1.8) $R^{h}\left(w,\overline{w},v,\overline{v}\right)\leq R^{g}\left(w,\overline{w},v,\overline{v}\right),$
and $R^{g}\left(w,\overline{w},v,\overline{v}\right)>0$, then $\omega_{h}\leq\omega_{g}$.
As an application, we obtain a rigid characterization for Hermitian metrics.
Corollary 1.4.
Let $g$ and $h$ be two Hermitian metrics on a compact complex manifold $M$. If for every $x\in M$ and every nonzero $v\in T_{x}^{1,0}M$, there exists some $w\in T_{x}^{1,0}M$ such that $R^{h}\left(w,\overline{w},v,\overline{v}\right)=R^{g}\left(w,\overline{w},v,\overline{v}\right)>0$, then $\omega_{h}=\omega_{g}$.
As another application of Theorem 1.3, we obtain the following diameter and volume comparison theorem by using RC-positivity.
Corollary 1.5.
Let $g$ and $h$ be two Hermitian metrics on a compact complex manifold $M$. If for every $x\in M$ and every nonzero $v\in T_{x}^{1,0}M$, there exists some $w\in T_{x}^{1,0}M$ such that $R^{g}\left(w,\overline{w},v,\overline{v}\right)>0$, and
(1.9) $R^{h}\left(w,\overline{w},v,\overline{v}\right)\leq R^{g}\left(w,\overline{w},v,\overline{v}\right),$
then $\mathrm{diam}(M,h)\leq\mathrm{diam}(M,g)$ and $\mathrm{Vol}(M,h)\leq\mathrm{Vol}(M,g)$.
The following special case is of particular interest.
Corollary 1.6.
Let $(M,g)$ be a compact Hermitian manifold with positive holomorphic sectional curvature. If $h$ is another Hermitian metric on $M$ satisfying
(1.10) $R^{h}(v,\overline{v},v,\overline{v})\leq R^{g}(v,\overline{v},v,\overline{v}),$
for any $v\in T^{1,0}M$, then $\omega_{h}\leq\omega_{g}$. In particular,
(1.11) $\mathrm{diam}(M,h)\leq\mathrm{diam}(M,g)\quad\mbox{ and }\quad\mathrm{Vol}(M,h)\leq\mathrm{Vol}(M,g).$
We also obtain similar comparison theorems for complex manfiolds with negative holomorphic sectional curvature.
Theorem 1.7.
Let $(M,h)$ be a compact Hermitian manifold with negative holomorphic sectional curvature. If $g$ is another Hermitian metric on $M$ satisfying
(1.12) $\mathrm{HSC}(h,v)\leq\mathrm{HSC}(g,v),$
for each nonzero $v\in T^{1,0}M$, then
(1.13) $\omega_{h}\leq\omega_{g}.$
In particular, we have $\mathrm{diam}(M,h)\leq\mathrm{diam}(M,g)$ and $\mathrm{Vol}(M,h)\leq\mathrm{Vol}(M,g)$.
The proofs of Theorem 1.3 and Theorem 1.7 are obtained in a general abstract vector bundle setting. Let’s fix the setup for readers’ convenience. Let $f:M\rightarrow N$ be a holomorphic map between complex manifolds and $g$, $h$ be Hermitian metrics on $M$ and $N$ respectively. Suppose that $(E_{1},G)\rightarrow M$ and $(E_{2},H)\rightarrow N$ are two Hermitian holomorphic vector bundles. If $\varphi:E_{1}\to E_{2}$ is a holomorphic bundle map, then there is an induced bundle map $\Phi:E_{1}\rightarrow f^{*}E_{2}$. Let $\widehat{E}_{2}=f^{*}E_{2}$ be the pullback bundle over $M$ and $\widehat{H}$ be the induced metric on $\widehat{E}_{2}$. The Chern curvature tensors of $(E_{1},G)$, $(E_{2},H)$ and $(\widehat{E}_{2},\widehat{H})$ are denoted by $R^{E_{1}}$, $R^{E_{2}}$ and $R^{\widehat{E}_{2}}$ respectively. Without loss of generality, we assume that $M$ is compact; $f:M\rightarrow N$ and $\varphi:E_{1}\rightarrow E_{2}$ are non-trivial.
1.A. Schwarz lemmas for RC-positive vector bundles
By using the previous setup, we obtain Schwarz lemmas for holomorphic bundle maps between abstract Hermitian vector bundles whose curvature tensors have positive bounds.
Theorem 1.8.
If there is a constant $\kappa\in{\mathbb{R}}$ such that: for every $x\in M$ and every non-zero $\xi\in(E_{1})_{x}$, there exists some $w\in T_{x}^{1,0}M$ satisfying
(1.14) $R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)>0$
and
(1.15) $R^{\widehat{E}_{2}}\left(w,\overline{w},\Phi(\xi),\overline{\Phi(\xi)}\right)\leq\kappa\cdot R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right),$
then $\kappa>0$ and
(1.16) $\varphi^{*}H\leq\kappa\cdot G.$
The proof of Theorem 1.8 relies on a maximum principle for abstract vector bundles with desired curvature positivity. There are many applications of Theorem 1.8. For instance, if $(E_{1},G)=\left(T^{1,0}M,g\right)$, $\left(E_{1},H\right)=\left(T^{1,0}N,h\right)$ and $\varphi=f_{*}$, we obtain:
Corollary 1.9.
Let $(M,g)$ be a compact Hermitian manifold and $(N,h)$ be a Hermitian manifold. If $f:M\rightarrow N$ is a non-constant holomorphic map such that for every $x\in M$ and every nonzero $v\in T_{x}^{1,0}M$, there exists some $w\in T_{x}^{1,0}M$ satisfying
(1.17) $R^{g}\left(w,\overline{w},v,\overline{v}\right)>0\quad\mbox{and}\quad R^{h}\left(f_{*}w,\overline{f_{*}w},f_{*}v,\overline{f_{*}v}\right)\leq\kappa\cdot R^{g}\left(w,\overline{w},v,\overline{v}\right),$
where $\kappa\in{\mathbb{R}}$ is a constant, then $\kappa>0$ and
(1.18) $f^{*}\omega_{h}\leq\kappa\omega_{g}.$
As a special case, one has
Corollary 1.10.
Let $(M,g)$ be a compact Hermitian manifold with positive holomorphic sectional curvature, and $(N,h)$ be a Hermitian manifold. If $f:M\to N$ is a non-trivial holomorphic map such that for every $x\in M$ and every $v\in T_{x}^{1,0}M$ one has
(1.19) $R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)\leq\kappa\cdot R^{g}\left(v,\overline{v},v,\overline{v}\right),$
where $\kappa\in{\mathbb{R}}$ is a constant, then $\kappa>0$ and
(1.20) $f^{*}\omega_{h}\leq\kappa\omega_{g}.$
It is clear that Theorem 1.3 follows from Corollary 1.9, and Corollary 1.6 follows from Corollary 1.10.
1.B. Schwarz lemmas for RC-negative vector bundles
By using similar methods as in the proof of Theorem 1.8, we can derive Schwarz lemmas for holomorphic maps between abstract Hermitian vector bundles with negative curvature bounds, which extend classical Schwarz lemmas in [Yau78], [CCL79], [Roy80] and etc..
Theorem 1.11.
Let $(M,g)$ be a compact Hermitian manifold, and $(N,h)$ be a Hermitian manifold with negative holomorphic sectional curvature. If $f:M\to N$ is a non-constant holomorphic map such that for every $x\in M$ and every $v\in T_{x}^{1,0}M$ with $f_{*}v\neq 0$, one has
(1.21) $\kappa\cdot\mathrm{HSC}\left(h,f_{*}v\right)\leq\mathrm{HSC}\left(g,v\right),$
where $\kappa\in{\mathbb{R}}$ is a constant, then $\kappa>0$ and
(1.22) $f^{*}\omega_{h}\leq\kappa\omega_{g}.$
Theorem 1.7 is a consequence of Theorem 1.11.
As another straightforward application of Theorem 1.11, it is easy to deduce the following classical version of Schwarz’s lemma for holomorphic maps between complex manifolds with constant negative holomorphic sectional curvature bounds (e.g. [CCL79], [Roy80], [Ni19], [Bro22] and etc.).
Corollary 1.12.
Let $(M,g)$ and $(N,h)$ be two Hermitian manifolds. Suppose that there exist two constants $\kappa>0$ and $\lambda\in{\mathbb{R}}$ such that
(1.23) $\mathrm{HSC}(M,g)\geq-\lambda\quad\mbox{and}\quad\mathrm{HSC}(N,h)\leq-\kappa.$
If $M$ is compact and $f:M\rightarrow N$ is a non-constant holomorphic map, then $\lambda>0$ and
(1.24) $f^{*}\omega_{h}\leq\frac{\lambda}{\kappa}\omega_{g}.$
1.C. Schwarz lemmas for constant positive curvature bounds
It is well-known that classical Schwarz lemmas are derived from Chern-Lu identities for holomorphic maps between complex manifolds. We obtain a version of Chern-Lu type identity for holomorphic bundle maps between Hermitian vector bundles. As applications, we establish the following Schwarz lemma for abstract vector bundles with positive curvature bounds.
Theorem 1.13.
If there exist two constants $\lambda>0$ and $\kappa\in{\mathbb{R}}$ such that
(1.25) $\mathrm{tr}_{\omega_{g}}R^{E_{1}}\geq\lambda\varphi^{*}H,$
and for all $y\in N$, $v\in T_{y}^{1,0}N$ and $s\in(E_{2})_{y}$,
(1.26) $R^{E_{2}}\left(v,\overline{v},s,\overline{s}\right)\leq\kappa|v|_{h}^{2}|s|_{H}^{2},$
then $\kappa>0$ and
(1.27) $\sup_{M}\mathrm{tr_{G}}\varphi^{*}H\leq\frac{\kappa r_{0}}{\lambda}\left(\sup_{M}\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\right)$
where $r_{0}$ is the maximal rank of the bundle map $\varphi:E_{1}\rightarrow E_{2}$.
If we set $(E_{1},G)=\left(T^{1,0}M,g\right)$, $\left(E_{2},H\right)=\left(T^{1,0}N,h\right)$ and $\varphi=f_{*}$ in Theorem 1.13, then one can see clearly that
$\lambda\leq\kappa r_{0}.$
This is a version of Yau’s Schwarz lemma for constant positive curvature bounds. Indeed, we also obtain a version of Royden’s Schwarz lemma.
Theorem 1.14.
Let $(M,\omega_{g})$ be a complete Kähler manifold of dimension $n$ and $(N,\omega_{h})$ be a Kähler manifold. If $f:M\to N$ is a non-trivial holomorphic map and
(1.28) $\mathrm{Ric}(M,\omega_{g})\geq\lambda f^{*}\omega_{h}\quad\mbox{and}\quad\mathrm{HSC}(N,\omega_{h})\leq\kappa,$
where $\lambda>0$ and $\kappa>0$ are two constants, then
(1.29) $\lambda\leq\frac{\left(r_{f}+1\right)\kappa}{2},$
where $r_{f}$ is the maximal complex rank of the holomorphic tangent map $df$.
The key ingredient in the proof of Theorem 1.14 is a combined curvature estimate in the refined Schwarz calculation, which is different from the classical one for manifolds with constant negative curvature bounds. There is an explanation on the difference between estimates (1.5) and (1.29). When $M$ is compact, $\mathrm{Ric}(M,\omega_{g})\geq\lambda f^{*}\omega_{h}$ can be implied by the condition $\mathrm{Ric}(M,\omega_{g})\geq\lambda_{1}\omega_{g}$ since one has $\omega_{g}\geq c_{0}f^{*}\omega_{h}$ for some $c_{0}>0$. This constant $c_{0}$ plays a similar role as $\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}$. We also obtain a rigid characterization in a slightly general setting:
Theorem 1.15.
Let $(M,\omega_{g})$ be a compact Hermitian manifold of dimension $n$ and $(N,\omega_{h})$ be a Kähler manifold. If $f:M\to N$ is a non-trivial holomorphic map and
(1.30) $\mathrm{Ric}^{\mathrm{(2)}}(M,\omega_{g})\geq\lambda f^{*}\omega_{h}\quad\mbox{and}\quad\mathrm{HSC}(N,\omega_{h})\leq\kappa,$
where $\lambda>0$ and $\kappa>0$ are two constants, then
(1.31) $\lambda\leq\frac{\left(n+1\right)\kappa}{2}.$
Moreover, if the identity in (1.31) holds, then $f^{*}\omega_{h}=c\omega_{g}$ for some $c>0$ and $(M,\omega_{g})$ is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},2c^{-1}\kappa^{-1}\omega_{\mathrm{FS}}\right)$.
The following version is also of particular interest.
Theorem 1.16.
Let $(M,\omega_{g})$ be a compact Hermitian manifold of dimension $n$, $(N,\omega_{h})$ be a Kähler manifold, and $f:M\to N$ be a non-constant holomorphic map. If there exist two constants $\lambda>0$ and $\kappa\in{\mathbb{R}}$ such that
(1.32) $\mathrm{Ric}^{\mathrm{(2)}}(\omega_{g})\geq\lambda\omega_{g}\quad\mbox{and}\quad R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)\leq\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h},$
for any $v\in T^{1,0}M$, then $\kappa>0$ and
(1.33) $\lambda\leq\frac{\left(n+1\right)\kappa}{2}.$
Moreover, if the identity in (1.33) holds, then $(M,\omega_{g})$ is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},2\kappa^{-1}\omega_{\mathrm{FS}}\right)$.
1.D. Rigidity of holomorphic maps between vector bundles
It is well-known that Schwarz lemmas can give Liouville type rigidity theorems for holomorphic maps. In this subsection, we present Liouville type rigidity theorems corresponding to Schwarz lemmas in Theorem 1.8.
We recall the concept of RC-positivity and the relationships between various classical notions of positivity for readers’ convenience. Let $(E,H)$ be a Hermitian holomorphic vector bundle over a complex manifold $M$. We say that $(E,H)$ is $k$-RC positive, if for every $x\in M$ and every non-zero $\xi\in E_{x}$, there exists a $k$-dimensional linear subspace $W$ of $T_{x}^{1,0}M$ such that for all non-zero $w\in W$,
$R^{E}\left(w,\overline{w},\xi,\overline{\xi}\right)>0.$
We can define $k$-RC non-positivity, $k$-RC non-negativity and $k$-RC negativity in similar ways. Suppose that $(M,\omega_{g})$ is a Hermitian manifold of dimension $n$. The following statements follow easily and we refer to [Yang18] for more details.
  1. (1)
    If $\mathrm{Ric}(\omega_{g})>0$ or $\mathrm{HSC}(\omega_{g})>0$, then $(T^{1,0}M,g)$ is $1$-RC positive;
  2. (2)
    If $\mathrm{Ric}(\omega_{g})<0$ or $\mathrm{HSC}(\omega_{g})<0$, then $(T^{1,0}M,g)$ is $1$-RC negative;
  3. (3)
    If $\mathrm{Ric}(\omega_{g})\leq 0$ or $\mathrm{HSC}(\omega_{g})\leq 0$, then $(T^{1,0}M,g)$ is $1$-RC non-positive;
  4. (4)
    $\mathrm{HBSC}(\omega_{g})>0$ if and only if $(T^{1,0}M,g)$ is $n$-RC positive;
  5. (5)
    $\mathrm{HBSC}(\omega_{g})<0$ if and only if $(T^{1,0}M,g)$ is $n$-RC negative;
  6. (6)
    $\mathrm{HBSC}(\omega_{g})\leq 0$ if and only if $(T^{1,0}M,g)$ is $n$-RC non-positive.
We obtain a type of Liouville rigidity theorem for vector bundles corresponding to Theorem 1.8.
Theorem 1.17.
Let $f:M\rightarrow N$ be a holomorphic map. Suppose that the Hermitian vector bundle $(E_{1},G)$ is $k$-RC positive, and $(E_{2},H)$ is $s$-RC non-positive. If
$k+s>\dim N,$
then any holomorphic bundle map $\varphi:E_{1}\rightarrow E_{2}$ must be trivial.
In particular, one has the following rigidity theorem for holomorphic maps between complex manifolds.
Corollary 1.18.
Let $(M,g)$ be a compact Hermitian manifold such that $(T^{1,0}M,g)$ is $k$-RC positive, and $(N,h)$ be a Hermitian manifold such that $(T^{1,0}N,h)$ is $s$-RC non-positive. If $k+s>\dim N$, then any holomorphic map $f:M\to N$ must be constant.
By using the relationships between $k$-RC positivity and classical curvature notions, it is easy to see that Corollary 1.18 generalizes many classical Liouville theorems. Moreover, Corollary 1.18 can give many new rigidity theorems. It is well-known that there are many projective manifolds $X$ with $c_{1}(X)\leq 0$ which can contain rational curves, i.e., there exist non-constant holomorphic maps $f:{\mathbb{C}}{\mathbb{P}}^{1}\rightarrow X$. For instances,
  1. (1)
    projective $K3$ surfaces $X$ and $c_{1}(X)=0$;
  2. (2)
    quintic surfaces $X$ in ${\mathbb{P}}^{3}$ and $c_{1}(X)<0$;
  3. (3)
    some projective Calabi-Yau manifolds.
Note that, as the domain manifold of holomorphic maps into $X$, ${\mathbb{C}}{\mathbb{P}}^{1}$ has positive curvature operator. As an application of Corollary 1.18, we obtain rigidity theorems for holomorphic maps into target manifolds with $c_{1}(X)\leq 0$.
Corollary 1.19.
Let $X$ be a compact Kähler manifold with $c_{1}(X)\leq 0$ and $\dim_{\mathbb{C}}X=\ell$. If $(M,g)$ is a compact Hermitian manifold such that $(T^{1,0}M,g)$ is $\ell$-RC positive, then any holomorphic map $f:M\rightarrow X$ must be a constant map.
There is a rigidity theorem corresponding to Theorem 2.2.
Theorem 1.20.
Let $f:M\rightarrow N$ be a holomorphic map between complex manifolds and $r$ be the minimal rank of the holomorphic tangent map $df$. Suppose that $(E_{1},G)$ is $k$-RC non-negative and $(E_{2},H)$ is $s$-RC negative. If
$k+s>\dim M+\dim N-r,$
then any holomorphic bundle map $\varphi:E_{1}\rightarrow E_{2}$ is trivial.
In particular, one has
Corollary 1.21.
Let $(M,g)$ be a compact Hermitian manifold such that $(T^{1,0}M,g)$ is $k$-RC non-negative, and $(N,h)$ be a Hermitian manifold such that $(T^{1,0}N,h)$ is $s$-RC negative. If $f:M\to N$ is a holomorphic map and $r$ is the minimal complex rank of the holomorphic tangent map $df$, then
$r\leq\dim M+\dim N-k-s.$
Remark 1.22.
Many of the stated results also hold on almost complex manifolds.
Acknowledgements. The second named author would like to thank Valentino Tosatti for some helpful comments.
2. Schwarz’s lemma and RC-positivity
In this section, we establish Schwarz lemmas for abstract vector bundles with RC-positive or RC-negative curvature bounds, and prove Theorem 1.8 and Theorem 1.11. Let’s recall Theorem 1.8.
Theorem 2.1.
Let $f:M\rightarrow N$ be a non-constant holomorphic map between two complex manifolds and $M$ be compact. Let $(E_{1},G)\rightarrow M$ and $(E_{2},H)\rightarrow N$ be two Hermitian holomorphic vector bundles and $\varphi:E_{1}\to E_{2}$ be a non-trivial holomorphic bundle map. Suppose that for every $x\in M$ and every non-zero $\xi\in(E_{1})_{x}$, there exists some $w\in T_{x}^{1,0}M$ such that
(2.1) $R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)>0$
and
(2.2) $R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)\leq\kappa\cdot R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right),$
where $\kappa\in{\mathbb{R}}$ is a constant, then $\kappa>0$ and
(2.3) $\varphi^{*}H\leq\kappa\cdot G.$
Proof.
We define a continuous function $\mu:M\to{\mathbb{R}}$ as follows
(2.4) $\mu(x)=\sup_{\xi\in(E_{1})_{x}\setminus\{0\}}\frac{|\varphi(\xi)|_{H}^{2}}{|\xi|_{G}^{2}}.$
We shall prove that $\sup_{M}\mu\leq\kappa$. This implies $\kappa>0$ and the estimate in (2.3).
Since $M$ is compact, $\mu$ attains its maximum at some point $x_{0}\in M$. Moreover, there exists a non-zero $\xi\in(E_{1})_{x_{0}}$ such that
$\mu(x_{0})=\frac{|\varphi(\xi)|_{H}^{2}}{|\xi|_{G}^{2}}.$
We extend $\xi$ to a local non-vanishing holomorphic section $s\in H^{0}(U,E_{1})$, where $U$ is an open neighborhood of $p\in M$, such that
(2.5) $s(x_{0})=\xi\quad\mbox{and}\quad(\nabla^{G}s)(x_{0})=0,$
where $\nabla^{G}$ is the Chern connection of $(E_{1},G)$. Note that $\varphi$ induces a holomorphic bundle map $\Phi=\varphi_{*}:E_{1}\to\widehat{E}_{2}=f^{*}E_{2}$ and
$\Phi(s)\in H^{0}\left(U,\widehat{E}_{2}\right).$
We define a function $F:U\to{\mathbb{R}}$ by
(2.6) $F(x)=\frac{|\Phi(s)|^{2}_{\widehat{H}}}{|s|_{G}^{2}}(x).$
Here $\widehat{H}$ is the induced metric on $\widehat{E}_{2}$. It is clear that
(2.7) $F(x_{0})=\sup_{U}F\quad\mbox{and}\quad F(x_{0})=\mu(x_{0}).$
Since $\mu(x_{0})>0$, $F$ is positive near $x_{0}$. Hence, at $x_{0}\in M$, for any $w\in T^{1,0}_{x_{0}}M$, one has
(2.8) $0\geq\left(\partial\overline{\partial}\log F\right)(w,\overline{w})=\left(\partial\overline{\partial}\log|\Phi(s)|^{2}_{\widehat{H}}\right)(w,\overline{w})-\left(\partial\overline{\partial}\log|s|_{G}^{2}\right)(w,\overline{w}).$
Note that $\Phi(s)(x_{0})=\varphi(\xi)$. Let $\widehat{\nabla}$ be the induced connection on $\left(\widehat{E}_{2},\widehat{H}\right)$. At point $x_{0}$, one has the estimate:
(2.9) (¯$\left(\partial\overline{\partial}\log|\Phi(s)|^{2}_{\widehat{H}}\right)(w,\overline{w})\geq-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
Indeed, a straightforward computation shows that
(¯$\left(\partial\overline{\partial}\log|\Phi(s)|^{2}_{\widehat{H}}\right)(w,\overline{w})=\frac{\left(\partial\overline{\partial}|\Phi(s)|^{2}_{\widehat{H}}\right)(w,\overline{w})}{|\varphi(\xi)|^{2}_{H}}-\frac{\left(\partial|\Phi(s)|^{2}_{\widehat{H}}\wedge\overline{\partial}|\Phi(s)|^{2}_{\widehat{H}}\right)(w,\overline{w})}{|\varphi(\xi)|^{4}_{H}}$
By using Bochner technique, the right hand side can be written as
|^$\frac{\left|\widehat{\nabla}_{w}(\Phi(s))\right|_{H}^{2}-R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}-\frac{\left|\left\langle\widehat{\nabla}_{w}(\Phi(s)),\varphi(\xi)\right\rangle_{H}\right|^{2}}{|\varphi(\xi)|^{4}_{H}}.$
On the other hand, we have the Cauchy-Schwarz inequality
(2.10) |^$\left|\left\langle\widehat{\nabla}_{w}(\Phi(s)),\varphi(\xi)\right\rangle_{H}\right|^{2}\leq\left|\widehat{\nabla}_{w}(\Phi(s))\right|_{H}^{2}\cdot|\varphi(\xi)|_{H}^{2}.$
Hence, we obtain (2.9). Similarly, at point $x_{0}$, one has
(¯$\left(\partial\overline{\partial}\log|s|^{2}_{G}\right)(w,\overline{w})=\frac{\left|\nabla^{G}_{w}s\right|_{G}^{2}-R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{\left|\left\langle\nabla^{G}_{w}s,\xi\right\rangle_{G}\right|^{2}}{|\xi|_{G}^{4}}.$
By using the fact that $(\nabla^{G}s)(x_{0})=0$, we obtain
(¯$\left(\partial\overline{\partial}\log|s|^{2}_{G}\right)(w,\overline{w})=-\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}.$
Therefore we get the inequality
(2.11) $0\geq\left(\partial\overline{\partial}\log F\right)(w,\overline{w})\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
In particular, if we take $w$ as in (2.1) and (2.2), then
(2.12) $0\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{\kappa R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\varphi(\xi)|^{2}_{H}}.$
By using (2.1), one obtains
(2.13) $\mu(x_{0})=F(x_{0})=\frac{|\varphi(\xi)|^{2}_{H}}{|\xi|_{G}^{2}}\leq\kappa,$
which completes the proof of Theorem 2.1. ∎
Proof of Theorem 1.11. By using a similar setup as in the proof of Theorem 2.1, we define
$\mu(x)=\sup_{x\in T_{x}^{1,0}M\setminus\{0\}}\frac{|f_{*}v|_{h}^{2}}{|v|_{g}^{2}}.$
Suppose that $\mu$ attains its maximum at some point $x_{0}\in M$. Moreover, there exists some non-zero $v\in T_{x_{0}}^{1,0}M$ such that
$\mu(x_{0})=\frac{|f_{*}v|_{h}^{2}}{|v|_{g}^{2}}.$
We extend $v$ to a local non-vanishing holomorphic vector field $V$ in an open neighborhood $U$ of $p\in M$, such that
(2.14) $V(x_{0})=v\quad\mbox{and}\quad(\nabla^{g}V)(x_{0})=0,$
where $\nabla^{g}$ is the Chern connection of $(T^{1,0}M,g)$. We define a function $F:U\to{\mathbb{R}}$ by
(2.15) $F(x)=\frac{|f_{*}V|^{2}_{h}}{|V|_{g}^{2}}(x).$
By using a similar argument as in the proof of Theorem 2.1, we obtain the following estimate at point $x_{0}\in M$: for $v=w\in T_{x_{0}}^{1,0}M$
(2.16) $0\geq\left(\partial\overline{\partial}\log F\right)(v,\overline{v})\geq\frac{R^{g}\left(v,\overline{v},v,\overline{v}\right)}{|v|_{g}^{2}}-\frac{R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)}{|f_{*}v|^{2}_{h}}.$
By the curvature condition (1.21), we obtain
(2.17) $0\geq\left(\kappa\cdot\frac{|v|^{2}_{g}}{|f_{*}v|_{h}^{2}}-1\right)\frac{R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)}{|f_{*}v|^{2}_{h}}.$
Since $f_{*}v\neq 0$, we have $R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)<0$, and so
(2.18) $\mu(p)=\frac{|f_{*}v|^{2}_{h}}{|v|^{2}_{g}}\leq\kappa.$
This is $f^{*}\omega_{h}\leq\kappa\cdot\omega_{g}$.∎
Theorem 2.2.
If there is a constant $\kappa\in{\mathbb{R}}$ such that: for every $x\in M$ and every non-zero $\xi\in(E_{1})_{x}$, there exists some $w\in T_{x}^{1,0}M$ satisfying
(2.19) $R^{\widehat{E}_{2}}\left(w,\overline{w},\Phi(\xi),\overline{\Phi(\xi)}\right)<0$
and
(2.20) $\kappa\cdot\frac{R^{\widehat{E}_{2}}\left(w,\overline{w},\Phi(\xi),\overline{\Phi(\xi)}\right)}{|w|^{2}_{g}\cdot|\Phi(\xi)|^{2}_{\widehat{H}}}\leq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|w|_{g}^{2}\cdot|\xi|^{2}_{G}},$
then
$\kappa\geq 1.$
Proof.
By using the same setup as in the proof of Theorem 2.1, we have the following estimate at point $x_{0}\in M$, i.e., for any $w\in T_{x_{0}}^{1,0}M$
(2.21) $0\geq\left(\partial\overline{\partial}\log F\right)(w,\overline{w})\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
By the curvature condition (2.20), we obtain that
$\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}\geq(\kappa-1)\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
Since $R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)<0$, we obtain $\kappa-1\geq 0$ and that is $\kappa\geq 1$. ∎
3. Schwarz’s lemma and non-negative second Chern-Ricci curvature
In this section, we establish Chern-Lu type identities for holomorphic bundle maps and prove Theorem 1.13. We recall the setup for readers’ convenience. Let $f:M\to N$ be a holomorphic map. Suppose that $(E_{1},G)\rightarrow M$ and $(E_{2},H)\rightarrow N$ are two Hermitian vector bundles. If $\varphi:E_{1}\to E_{2}$ is a holomorphic bundle map, then there is an induced bundle map $\Phi:E_{1}\rightarrow f^{*}E_{2}$. Let $\widehat{E}_{2}=f^{*}E_{2}$ and $\widehat{H}$ be the induced metric on $\widehat{E}_{2}$. For simplicity, we set $E=E_{1}^{*}\otimes\widehat{E}_{2}$ and there is an induced metric $G^{*}\otimes\widehat{H}$ on $E$. Since $\varphi$ is a holomorphic bundle map, $\Phi$ is a holomorphic section of $E$, i.e. $\Phi\in H^{0}(M,E)$. We have the following Bochner formula
(3.1) ¯$\partial\overline{\partial}|\Phi|^{2}=\left\{\nabla\Phi,\nabla\Phi\right\}-\left\{R^{E}(\Phi),\Phi\right\},$
where $R^{E}$ is the Chern curvature of $(E,G^{*}\otimes\widehat{H})$.
Suppose that $E_{1}$ and $E_{2}$ have complex ranks $r_{1}$ and $r_{2}$ respectively. Let $\{e_{a}\}_{1\leq a\leq r_{1}}$ and $\{\widetilde{e}_{A}\}_{1\leq A\leq r_{2}}$ be local holomorphic frames of $E_{1}$ and $E_{2}$ respectively. The bundle map $\varphi$ is given by
(3.2) $\varphi(e_{a})=\sum_{A}\varphi_{a}^{A}\cdot\widetilde{e}_{A},$
where $\varphi_{a}^{A}$ are local holomorphic functions on $M$. If we set $\widehat{e}_{A}=f^{*}\widetilde{e}_{A}$, then
(3.3) $\Phi(e_{a})=\sum_{A}\varphi_{a}^{A}\cdot\widehat{e}_{A}.$
Let $\{z^{i}\}$ and $\{w^{\alpha}\}$ be local holomorphic coordinates on $M$ and $N$ respectively. The holomorphic tangent map $df$ is represented by $f_{i}^{\alpha}=\frac{\partial f^{\alpha}}{\partial z^{i}}$. It is easy to see that the Chern curvature of $\left(\widehat{E}_{2},\widehat{H}\right)$ is given by
(3.4) $R^{E_{2}}_{\alpha\overline{\beta}A\overline{B}}f_{i}^{\alpha}\overline{f_{j}^{\beta}}dz^{i}\wedge d\overline{z}^{j}\otimes\widehat{e}^{A}\otimes\overline{\widehat{e}^{B}}.$
By using the curvature formula
$R^{E}=R^{E_{1}^{*}}\otimes\mathrm{id}_{\widehat{E}_{2}}+\mathrm{id}_{E_{1}^{*}}\otimes R^{\widehat{E}_{2}},$
we obtain Chern-Lu type identities for holomorphic bundle maps.
Lemma 3.1.
The local expression of formula (3.1) is
(3.5) ¯$\partial\overline{\partial}|\Phi|^{2}=\left\{\nabla\Phi,\nabla\Phi\right\}+\left[R^{E_{1}}_{i\overline{j}a\overline{b}}G^{a\overline{d}}G^{c\overline{b}}\varphi_{c}^{A}\overline{\varphi_{d}^{B}}H_{A\overline{B}}-R^{E_{2}}_{\alpha\overline{\beta}A\overline{B}}f_{i}^{\alpha}\overline{f_{j}^{\beta}}\varphi_{a}^{A}\overline{\varphi_{b}^{B}}G^{a\overline{b}}\right]dz^{i}\wedge d\overline{z}^{j}.$
Moreover,
(3.6) $\Box_{g}|\Phi|^{2}=|\nabla\Phi|^{2}+g^{i\overline{j}}R^{E_{1}}_{i\overline{j}a\overline{b}}G^{a\overline{d}}G^{c\overline{b}}\varphi_{c}^{A}\overline{\varphi_{d}^{B}}H_{A\overline{B}}-R^{E_{2}}_{\alpha\overline{\beta}A\overline{B}}g^{i\overline{j}}f_{i}^{\alpha}\overline{f_{j}^{\beta}}\varphi_{a}^{A}\overline{\varphi_{b}^{B}}G^{a\overline{b}},$
where $\Box_{g}|\Phi|^{2}=\mathrm{tr}_{\omega_{g}}(\sqrt{-1}\partial\overline{\partial}|\Phi|^{2})$.
As an application of Lemma 3.1, we obtain Theorem 1.13. That is:
Theorem 3.2.
Let $f:(M,g)\rightarrow(N,h)$ be a non-constant holomorphic map between two Hermitian manifolds and $M$ be compact. Suppose that $(E_{1},G)\rightarrow M$ and $(E_{2},H)\rightarrow N$ are two Hermitian vector bundles, and $\varphi:E_{1}\to E_{2}$ is a non-trivial holomorphic bundle map. Suppose that there exist two constants $\lambda>0$ and $\kappa>0$ such that
(3.7) $\mathrm{tr}_{\omega_{g}}R^{E_{1}}\geq\lambda\varphi^{*}H,$
and for all $y\in N$, $v\in T_{y}^{1,0}N$ and $s\in(E_{2})_{y}$, there holds
(3.8) $R^{E_{2}}\left(v,\overline{v},s,\overline{s}\right)\leq\kappa|v|_{h}^{2}|s|_{H}^{2}.$
Then one has
(3.9) $\sup_{M}\mathrm{tr_{G}}\varphi^{*}H\leq\frac{\kappa r_{0}}{\lambda}\left(\sup_{M}\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\right)$
where $r_{0}$ is the maximal rank of the bundle map $\varphi:E_{1}\rightarrow E_{2}$.
Proof.
At a fixed point $x\in M$, there exist local holomorphic coordinates $\{z^{i}\}$ and $\{w^{\alpha}\}$ around $x\in M$ and $f(x)\in N$ respectively such that
$g_{i\overline{j}}(x)=\delta_{ij},\quad h_{\alpha\overline{\beta}}(f(x))=\delta_{\alpha\beta}\quad\mbox{and}\quad f_{i}^{\alpha}(x)=\lambda_{i}\delta_{i}^{\alpha}$
with $\lambda_{1}\geq\cdots\geq\lambda_{r}>\lambda_{r+1}=\cdots=\lambda_{n}=0$, where $r$ is the complex rank of $\left[f_{i}^{\alpha}\right]$. Similarly, there exist local holomorphic frames $\{e_{a}\}$ and $\{\widetilde{e}_{A}\}$ of $E_{1}$ and $E_{2}$ respectively such that
$G_{a\overline{b}}(x)=\delta_{ab},\quad H_{A\overline{B}}(f(x))=\delta_{AB}\quad\mbox{and}\quad\varphi_{a}^{A}(x)=\Lambda_{a}\delta_{a}^{A}$
with $\Lambda_{1}\geq\cdots\geq\Lambda_{r^{\prime}}>\Lambda_{r^{\prime}+1}=\cdots=\Lambda_{r_{1}}=0$, where $r^{\prime}$ is the rank of $\left[\varphi_{a}^{A}\right]$. It is easy to see that
$|\Phi|^{2}=\mathrm{tr_{G}}\varphi^{*}H=G^{a\overline{b}}H_{A\overline{B}}\varphi^{A}_{a}\overline{\varphi^{B}_{b}}=\sum\Lambda_{a}^{2},\ \ \ |df|^{2}=\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}=\sum\lambda_{i}^{2}.$
By formulas (3.6), (3.7) and $r^{\prime}\leq\min\{r_{1},r_{2}\}$, one has
(3.10) $g^{i\overline{j}}R^{E_{1}}_{i\overline{j}a\overline{b}}G^{a\overline{d}}G^{c\overline{b}}\varphi_{c}^{A}\overline{\varphi_{d}^{B}}H_{A\overline{B}}=\sum_{a=1}^{r_{1}}\sum_{A=1}^{r_{2}}\mathrm{Ric}^{(2)}_{a\overline{a}}\Lambda_{a}^{2}\delta_{a}^{A}\geq\lambda\sum_{a=1}^{r_{1}}\Lambda_{a}^{4}.$
Moreover, since $r_{0}\sum\Lambda_{a}^{4}\geq\left(\sum\Lambda_{a}^{2}\right)^{2}$, we obtain
(3.11) $g^{i\overline{j}}R^{E_{1}}_{i\overline{j}a\overline{b}}G^{a\overline{d}}G^{c\overline{b}}\varphi_{c}^{A}\overline{\varphi_{d}^{B}}H_{A\overline{B}}\geq\frac{\lambda}{r_{0}}|\Phi|^{4}.$
One the other hand, by formula (3.8), one has
$-R^{E_{2}}_{\alpha\overline{\beta}A\overline{B}}g^{i\overline{j}}f_{i}^{\alpha}\overline{f_{j}^{\beta}}\varphi_{a}^{A}\overline{\varphi_{b}^{B}}G^{a\overline{b}}$ $=$ $-\sum_{i=1}^{n}\sum_{\alpha=1}^{m}\sum_{a=1}^{r_{1}}\sum_{A=1}^{r_{2}}R^{E_{2}}_{\alpha\overline{\alpha}A\overline{A}}\lambda_{i}^{2}\delta_{i}^{\alpha}\Lambda_{a}^{2}\delta_{a}^{A}$
$\geq$ $-\kappa\sum_{i=1}^{n}\sum_{\alpha=1}^{m}\sum_{a=1}^{r_{1}}\sum_{A=1}^{r_{2}}\lambda_{i}^{2}\delta_{i}^{\alpha}\Lambda_{a}^{2}\delta_{a}^{A}$
$=$ $-\kappa\sum_{i=1}^{n}\lambda_{i}^{2}\sum_{a=1}^{r_{1}}\Lambda_{a}^{2}.$
Here we use facts that $r\leq\min\{m,n\}$ and $r^{\prime}\leq\min\{r_{1},r_{2}\}$. By (3.6), we obtain
(3.12) $\Box_{g}|\Phi|^{2}\geq\frac{\lambda}{r_{0}}|\Phi|^{4}-\kappa|\Phi|^{2}|df|^{2}.$
Suppose $|\Phi|^{2}$ attains its maximum at point $p\in M$. By using maximum principle,
$\sup_{M}\mathrm{tr_{G}}\varphi^{*}H=\left(\mathrm{tr_{G}}\varphi^{*}H\right)(p)\leq\frac{\kappa r_{0}}{\lambda}\left(\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\right)(p)\leq\frac{\kappa r_{0}}{\lambda}\left(\sup_{M}\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\right).$
This completes the proof. ∎
In particular, one has
Corollary 3.3.
Let $(E_{1},G)$ and $(E_{2},H)$ be two Hermitian vector bundles over a compact Hermitian manifold $(M,g)$, and $\varphi:E_{1}\to E_{2}$ be a non-trivial holomorphic bundle map. Suppose that there exist two constants $\lambda>0$ and $\kappa>0$ such that
(3.13) $\mathrm{tr}_{\omega_{g}}R^{E_{1}}\geq\lambda\varphi^{*}H,$
and for all $x\in M$, $v\in T_{x}^{1,0}M$ and $s\in(E_{2})_{x}$, there holds
(3.14) $R^{E_{2}}\left(v,\overline{v},s,\overline{s}\right)\leq\kappa|v|_{g}^{2}|s|_{H}^{2}.$
Then one has
(3.15) $\sup_{M}\mathrm{tr_{G}}\varphi^{*}H\leq\frac{n\kappa r_{0}}{\lambda}$
where $\dim M=n$ and $r_{0}$ is the maximal rank of the bundle map $\varphi:E_{1}\rightarrow E_{2}$.
By a similar method as in the proof of Theorem 3.2, one can show:
Theorem 3.4.
Let $f:(M,g)\rightarrow(N,h)$ be a non-constant holomorphic map and $M$ be compact. Suppose that $(E_{1},G)\rightarrow M$ and $(E_{2},H)\rightarrow N$ are two Hermitian vector bundles, and $\varphi:E_{1}\to E_{2}$ is a non-trivial holomorphic bundle map. Suppose that there exist two constants $\lambda\in{\mathbb{R}}$ and $\kappa\in{\mathbb{R}}$ such that
(3.16) $\mathrm{tr}_{\omega_{g}}R^{E_{1}}\geq\lambda G,$
and for all $x\in M$, $v\in T_{x}^{1,0}M$ and $s\in(f^{*}E_{2})_{x}$,
(3.17) $R^{f^{*}E_{2}}\left(v,\overline{v},s,\overline{s}\right)\leq\kappa|v|_{g}^{2}|s|_{H}^{2}.$
Then one has
(3.18) $\lambda\leq\kappa\cdot\dim M.$
4. Schwarz’s lemma for positive curvature bounds and rigid characterizations
In this section, we prove Theorem 1.14, Theorem 1.15 and Theorem 1.16. We begin with the well-known Schwarz calculation (e.g. [Che68], [Lu68], [Yau78], [Yang21]), which is also a special case of Lemma 3.1.
Lemma 4.1.
Let $f:(M,\omega_{g})\rightarrow(N,\omega_{h})$ be a holomorphic map between Hermitian manifolds. Then in local holomorphic coordinates $\{z^{i}\}$ and $\{w^{\alpha}\}$ on $M$ and $N$, respectively, we have
(4.1) $\Box_{g}u=|\nabla df|^{2}+\left(g^{i\overline{j}}R^{g}_{i\overline{j}k\overline{\ell}}\right)g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}-R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\left(g^{i\overline{j}}f^{\alpha}_{i}\overline{f^{\beta}_{j}}\right)\left(g^{p\overline{q}}f^{\gamma}_{p}\overline{f^{\delta}_{q}}\right),$
where $u=\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}$, $f^{\alpha}_{i}=\frac{\partial f^{\alpha}}{\partial z^{i}}$, $\nabla$ is the induced Chern connection on $T^{*1,0}M\otimes f^{*}T^{1,0}N$, and $\Box_{g}u=\mathrm{tr}_{\omega_{g}}(\sqrt{-1}\partial\overline{\partial}u)$ is the complex Laplacian of $u$.
Proof of Theorem 1.14. Without loss of generality, we can assume $\lambda>\frac{1}{2}\kappa$. In this case, we claim that
(4.2) $\Box_{g}u\geq\frac{1}{r_{f}}\left(\lambda-\frac{(r_{f}+1)\kappa}{2}\right)u^{2}$
where $u=\mbox{tr}_{\omega_{g}}f^{*}\omega_{h}$ and $r_{f}$ is the maximum rank of $df$.
At a fixed point $x\in M$, there exist local holomorphic coordinates $\{z^{i}\}$ and $\{w^{\alpha}\}$ around $x\in M$ and $f(x)\in N$ respectively such that
$g_{i\overline{j}}(x)=\delta_{ij},\quad h_{\alpha\overline{\beta}}(f(x))=\delta_{\alpha\beta}\quad\mbox{and}\quad f_{i}^{\alpha}(x)=\lambda_{i}\delta_{i}^{\alpha}$
with $\lambda_{1}\geq\cdots\geq\lambda_{r}>\lambda_{r+1}=\cdots=\lambda_{n}=0$, where $r$ is the complex rank of $\left[f_{i}^{\alpha}\right]$. Clearly,
$u=\mbox{tr}_{\omega_{g}}f^{*}\omega_{h}=\sum_{i=1}^{n}\lambda_{i}^{2},$
and by using the curvature condition (1.28), one has
(4.3) $\left(g^{i\overline{j}}R^{g}_{i\overline{j}k\overline{\ell}}\right)g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}\geq\lambda\left(h_{\gamma\overline{\delta}}f_{k}^{\gamma}\overline{f_{\ell}^{\delta}}\right)g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}=\lambda\sum_{i=1}^{n}\lambda_{i}^{4}.$
By using the Kähler symmetry of $R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}$, one obtains
$\frac{2}{n(n+1)}R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\left(g^{i\overline{j}}f^{\alpha}_{i}\overline{f^{\beta}_{j}}\right)\left(g^{p\overline{q}}f^{\gamma}_{p}\overline{f^{\delta}_{q}}\right)$ $=$ $R^{h}_{i\overline{j}k\overline{\ell}}\lambda_{i}\lambda_{j}\lambda_{k}\lambda_{\ell}\frac{\delta_{ij}\delta_{k\ell}+\delta_{i\ell}\delta_{kj}}{n(n+1)}$
$=$ $\int_{{\mathbb{P}}^{n-1}}\frac{R^{h}_{i\overline{j}k\overline{\ell}}\lambda_{i}\xi^{i}\cdot\overline{\lambda_{j}\xi^{j}}\cdot\lambda_{k}\xi^{k}\cdot\overline{\lambda_{\ell}\xi^{\ell}}}{|\xi|^{4}}\omega_{\mathrm{FS}}^{n-1}$
where $R^{h}_{i\overline{j}k\overline{\ell}}=R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\delta_{i}^{\alpha}\delta_{j}^{\beta}\delta_{k}^{\gamma}\delta_{\ell}^{\delta}$. By utilizing the curvature condition in (1.28), we get
(4.4) $R^{h}_{i\overline{j}k\overline{\ell}}\lambda_{i}\xi^{i}\cdot\overline{\lambda_{j}\xi^{j}}\cdot\lambda_{k}\xi^{k}\cdot\overline{\lambda_{\ell}\xi^{\ell}}=R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}V^{\alpha}\overline{V^{\beta}}V^{\gamma}\overline{V^{\delta}}\leq\kappa|V|^{4}_{h}$
where $V^{\alpha}=\sum_{i=1}^{n}\delta_{i}^{\alpha}\lambda_{i}\xi^{i}$. It is easy to see that at the fixed point $f(x)$,
(4.5) $|V|^{2}_{h}=\sum_{\alpha}|V^{\alpha}|^{2}=\sum_{i=1}^{n}\lambda_{i}^{2}|\xi^{i}|^{2}.$
Moreover, we obtain
$\int_{{\mathbb{P}}^{n-1}}\frac{\sum_{i=1}^{n}\lambda_{i}^{2}|\xi^{i}|^{2}\cdot\sum_{k=1}^{n}\lambda_{k}^{2}|\xi_{k}|^{2}}{|\xi|^{4}}\omega_{\mathrm{FS}}^{n-1}$ $=$ $\int_{{\mathbb{P}}^{n-1}}\sum_{i,k=1}^{n}\lambda_{i}^{2}\lambda_{k}^{2}\frac{\xi^{i}\overline{\xi^{i}}\xi^{k}\overline{\xi^{k}}}{|\xi|^{4}}\omega_{\mathrm{FS}}^{n-1}$
$=$ $\frac{1}{n(n+1)}\left[\left(\sum_{i=1}^{n}\lambda_{i}^{2}\right)^{2}+\sum_{i=1}^{n}\lambda_{i}^{4}\right].$
Hence, we obtain
(4.6) $R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\left(g^{i\overline{j}}f^{\alpha}_{i}\overline{f^{\beta}_{j}}\right)\left(g^{p\overline{q}}f^{\gamma}_{p}\overline{f^{\delta}_{q}}\right)\leq\frac{\kappa}{2}\left(u^{2}+\sum_{i=1}^{n}\lambda_{i}^{4}\right).$
By using (4.3) and (4.6), (4.1) is reduced to
(4.7) $\Box_{g}u\geq\left(\lambda-\frac{\kappa}{2}\right)\sum_{i=1}^{n}\lambda_{i}^{4}-\frac{\kappa}{2}u^{2}.$
The Cauchy-Schwarz inequality gives $r_{f}\sum_{i=1}^{n}\lambda_{i}^{4}\geq u^{2}$. Since $\lambda>\frac{\kappa}{2}$, we get
(4.8) $\Box_{g}u\geq\frac{1}{r_{f}}\left(\lambda-\frac{(r_{f}+1)\kappa}{2}\right)u^{2},$
and this is (4.2).
If $\sup_{M}u<+\infty$, by Yau’s maximum principle, there is a sequence $\{p_{k}\}$ such that
(4.9) $\lim_{k\to\infty}u(p_{k})=\sup_{M}u\quad\mbox{and}\quad\limsup_{k\to\infty}\Box_{g}u(p_{k})\leq 0.$
Therefore,
$0\geq\limsup_{k\to\infty}\Box_{g}u(p_{k})\geq\lim_{k\to\infty}\frac{1}{r_{f}}\left(\lambda-\frac{r_{f}+1}{2}\kappa\right)u(p_{k})^{2}=\frac{1}{r_{f}}\left(\lambda-\frac{r_{f}+1}{2}\kappa\right)(\sup_{M}u)^{2}.$
This implies $\lambda\leq(r_{f}+1)\kappa/2$.
If $\sup_{M}u=+\infty$, we fix some $\varepsilon>0$ and consider $v=1/\sqrt{u+\varepsilon}$. One can show
$\Box_{g}v=-\frac{1}{2}v^{3}\Box_{g}u+\frac{3}{v}|\partial v|_{g}^{2}.$
By Yau’s maximum principle again, there exists a sequence $\{p_{k}\}$ such that
(4.10) $\lim_{k\to\infty}v(p_{k})=0,\quad\lim_{k\to\infty}|\partial v|_{g}(p_{k})=0\quad\mbox{and}\quad\liminf_{k\to\infty}\Box_{g}v(p_{k})\geq 0.$
Therefore,
$0\geq\limsup_{k\to\infty}\left(6|\partial v|_{g}^{2}-2v\Box_{g}v\right)(p_{k})$ $=$ $\limsup_{k\to\infty}\left(v^{4}\Box_{g}u\right)(p_{k})$
$\geq$ $\limsup_{k\to\infty}\left(\frac{1}{r_{f}}\left(\lambda-\frac{r_{f}+1}{2}\kappa\right)v^{4}u^{2}\right)(p_{k}).$
Since $\lim_{k\to\infty}v(p_{k})=0$, one has $\lim_{k\to\infty}u(p_{k})=+\infty$, and so
(4.11) $\lim_{k\to\infty}(v^{4}u^{2})(p_{k})=\lim_{k\to\infty}\frac{u^{2}}{(u+\varepsilon)^{2}}(p_{k})=1.$
This implies $\lambda\leq(r_{f}+1)\kappa/2$.∎
We refer to [LY17] for the definition of the second Chern-Ricci curvature $\mathrm{Ric}^{(2)}(\omega_{g})$ and related computations.
Proof of Theorem 1.15. Without loss of generality, we can assume $\lambda>\frac{1}{2}\kappa.$ By similar arguments as in the proof of Theorem 1.14 one has
(4.12) $\Box_{g}u\geq\frac{1}{r_{f}}\left(\lambda-\frac{(r_{f}+1)\kappa}{2}\right)u^{2}\geq\frac{1}{n}\left(\lambda-\frac{(n+1)\kappa}{2}\right)u^{2}$
where $u=\mbox{tr}_{\omega_{g}}f^{*}\omega_{h}$. Hence, $\lambda\leq\frac{(n+1)\kappa}{2}$.
Suppose $\lambda=(n+1)\kappa/2$. Then we have $\Box_{g}u\geq 0$. By using maximum principle again, we know $u$ is a constant. In this case, one can see clearly from (4.7) and (4.8) that $r_{f}=n$ and all $\lambda_{i}$ are the same, and so
$f^{*}\omega_{h}=c(p)\omega_{g}.$
Since $u=\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}$ is constant, $c(p)=c$ and $c>0$. This implies $cd\omega_{g}=f^{*}(d\omega_{h})=0$, and so $(M,\omega_{g})$ is Kähler. On the other hand, the inequality (4.6) is actually an identity, i.e., for any $v\in T^{1,0}M$
$R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)=\kappa|f_{*}v|^{4}_{h}.$
By (4.1), one has $\nabla df=0$. We claim that $g$ has constant holomorphic sectional curvature $c\kappa$, i.e., for any $p\in M$ and any non-zero $v\in T_{p}^{1,0}M$
(4.13) $\frac{R^{g}(v,\overline{v},v,\overline{v})}{|v|_{g}^{4}}=c\cdot\frac{R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)}{|f_{*}v|^{4}_{h}}=c\kappa.$
Indeed, for the given $v$, we extend it to a local non-vanishing holomorphic vector field $X$ over a neighborhood $U$ of $p$. A straightforward computation shows that at point $p$,
(4.14) (¯$\left(\partial\overline{\partial}\log|f_{*}X|^{2}_{h}\right)(v,\overline{v})=-\frac{R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)}{|f_{*}v|^{2}_{h}}+\frac{\left|\widehat{\nabla}_{v}(f_{*}X)\right|_{h}^{2}}{|f_{*}v|^{2}_{h}}-\frac{\left|\left\langle\widehat{\nabla}_{v}(f_{*}X),f_{*}v\right\rangle_{h}\right|^{2}}{|f_{*}v|^{4}_{h}}$
where $\widehat{\nabla}$ is the induced connection on $f^{*}T^{1,0}N$. Since $\nabla df=0$, we have
^$\widehat{\nabla}_{v}\left(f_{*}X\right)=f_{*}(\nabla_{v}X).$
Moreover, since $f^{*}\omega_{h}=c\omega_{g}$, we obtain that
(4.15) (¯$\left(\partial\overline{\partial}\log|f_{*}X|^{2}_{h}\right)(v,\overline{v})=-\frac{R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)}{|f_{*}v|^{2}_{h}}+\frac{\left|\nabla_{v}X\right|_{g}^{2}}{|v|^{2}_{g}}-\frac{\left|\left\langle\nabla_{v}X,v\right\rangle_{g}\right|^{2}}{|v|^{4}_{g}}.$
On the other hand,
(4.16) (¯$\left(\partial\overline{\partial}\log|X|^{2}_{g}\right)(v,\overline{v})=-\frac{R^{g}\left(v,\overline{v},v,\overline{v}\right)}{|v|^{2}_{g}}+\frac{\left|\nabla_{v}X\right|_{g}^{2}}{|v|^{2}_{g}}-\frac{\left|\left\langle\nabla_{v}X,v\right\rangle_{g}\right|^{2}}{|v|^{4}_{g}}.$
Since $|f_{*}X|^{2}_{h}=c|X|^{2}_{g}$, we obtain (4.13). Therefore, $(M,\omega_{g})$ is a compact Kähler manifold with positive constant holomorphic sectional curvature $c\kappa$, and so it is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},\frac{2}{c\kappa}\omega_{\mathrm{FS}}\right)$. ∎
Proof of Theorem 1.16. We claim that
(4.17) $\Box_{g}u\geq\left(\lambda-\frac{r_{f}+1}{2}\kappa\right)u\geq\left(\lambda-\frac{n+1}{2}\kappa\right)u,$
where $u=\mbox{tr}_{\omega_{g}}f^{*}\omega_{h}$ and $r_{f}$ is the maximum rank of $df$. Indeed, by the curvature condition in (1.32), we obtain
(4.18) $\left(g^{i\overline{j}}R^{g}_{i\overline{j}k\overline{\ell}}\right)g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}\geq\lambda g_{k\overline{\ell}}g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}=\lambda u.$
By using similar computations as in the proof of (4.6), one has
(4.19) $R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\left(g^{i\overline{j}}f^{\alpha}_{i}\overline{f^{\beta}_{j}}\right)\left(g^{p\overline{q}}f^{\gamma}_{p}\overline{f^{\delta}_{q}}\right)\leq\frac{\left(r_{f}+1\right)\kappa}{2}u.$
By formulas (4.18) and (4.19), we obtain (4.17). By using maximum principle, one can see clearly that $\lambda\leq\frac{n+1}{2}\kappa$.
Suppose $\lambda=(n+1)\kappa/2$. Then we have $\Box_{g}u\geq 0$ and so $u=C$ for some constant $C>0$. By (4.17), one obtains $r_{f}=n$. Moreover, for any $v\in T^{1,0}M$ one has
(4.20) $\mathrm{Ric}^{\mathrm{(2)}}(\omega_{g})=\lambda g\quad\mbox{and}\quad R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)=\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h}.$
Let $E=f^{*}T^{1,0}N$ be the pullback bundle, and $\widehat{\nabla}$ be the induced connection on $E$. By (4.1), one has $\nabla df\equiv 0$. This implies that for any vector field $X$ and $s\in\Gamma(M,T^{1,0}M)$
$f_{*}(\nabla_{X}^{g}s)=\widehat{\nabla}_{X}(f_{*}s),$
where $\nabla^{g}$ is the Chern connection of $(T^{1,0}M,g)$. In particular, for any $v,w\in T^{1,0}M$
(4.21) $f_{*}\left(R^{g}(v,\overline{v})w\right)=R^{E}(v,\overline{v})(f_{*}w).$
For simplicity, we write $H(v)=R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)$. At a fixed point $p$, let $\{e_{i}\}$ be a basis of $T_{p}^{1,0}M$ with $g(e_{i},e_{j})=\delta_{ij}$. A straightforward computation shows
$\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}H(e_{i}+te_{j})+\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}H\left(e_{i}+t\sqrt{-1}e_{j}\right)=16R^{h}\left(f_{*}e_{i},\overline{f_{*}e_{i}},f_{*}e_{j},\overline{f_{*}e_{j}}\right).$
On the other hand, since $H(v)=\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h}$ and $g(e_{i},e_{j})=\delta_{ij}$, one has
$\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}H(e_{i}+te_{j})+\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}H\left(e_{i}+t\sqrt{-1}e_{j}\right)=\begin{cases}4\kappa\left(\left|f_{*}e_{i}\right|_{h}^{2}+\left|f_{*}e_{j}\right|_{h}^{2}\right)&\text{if $i\neq j$,}\\ 16\kappa\left|f_{*}e_{i}\right|_{h}^{2}&\text{if $i=j$.}\end{cases}$
Therefore,
$R^{h}\left(f_{*}e_{i},\overline{f_{*}e_{i}},f_{*}e_{j},\overline{f_{*}e_{j}}\right)=\begin{cases}\frac{\kappa}{4}\left(\left|f_{*}e_{i}\right|_{h}^{2}+\left|f_{*}e_{j}\right|_{h}^{2}\right)&\text{if $i\neq j$},\\ \kappa\left|f_{*}e_{i}\right|_{h}^{2}&\text{if $i=j$.}\end{cases}$
Since $\sum_{i=1}^{n}\left|f_{*}e_{i}\right|_{h}^{2}=\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}=C$, we have
(4.22) $\sum_{i=1}^{n}R^{h}\left(f_{*}e_{i},\overline{f_{*}e_{i}},f_{*}e_{j},\overline{f_{*}e_{j}}\right)=\frac{C\kappa}{4}+\frac{(n+2)\kappa}{4}\left|f_{*}e_{j}\right|_{h}^{2}.$
On the other hand, by using (4.20) and (4.21), one has
$\sum_{i=1}^{n}R^{h}\left(f_{*}e_{i},\overline{f_{*}e_{i}},f_{*}e_{j},\overline{f_{*}e_{j}}\right)$
$=$ $\sum_{i=1}^{n}R^{E}\left(e_{i},\overline{e_{i}},f_{*}e_{j},\overline{f_{*}e_{j}}\right)=h\left(\sum_{i=1}^{n}R^{E}(e_{i},\overline{e_{i}})(f_{*}e_{j}),f_{*}e_{j}\right)$
$=$ $h\left(f_{*}\left(\sum_{i=1}^{n}R^{g}(e_{i},\overline{e_{i}})(e_{j})\right),f_{*}e_{j}\right)=h\left(\lambda f_{*}e_{j},f_{*}e_{j}\right)=\frac{(n+1)\kappa}{2}\left|f_{*}e_{j}\right|_{h}^{2}.$
Hence, $|f_{*}e_{j}|_{h}^{2}=\frac{C}{n}$ each $j$, and so $f^{*}\omega_{h}=\frac{C}{n}\omega_{g}$. By using similar arguments as in the proof of Theorem 1.15, we deduce that $g$ is a Kähler metric with constant holomorphic sectional curvature $\kappa$, and so $(M,\omega_{g})$ is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},2\kappa^{-1}\mathrm{\omega_{\mathrm{FS}}}\right)$. ∎
The constants $\lambda$ and $\kappa$ in Theorem 1.16 can also be negative.
Theorem 4.2.
Let $(M,\omega_{g})$ be a compact Hermitian manifold of dimension $n$, $(N,\omega_{h})$ be a Kähler manifold, and $f:M\to N$ be a non-constant holomorphic map. If there exist two constants $\lambda\in{\mathbb{R}}$ and $\kappa>0$ such that
(4.23) $\mathrm{Ric}^{\mathrm{(2)}}(\omega_{g})\geq-\lambda\omega_{g}\quad\mbox{and}\quad R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)\leq-\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h},$
for any $v\in T^{1,0}M$, then $\lambda>0$ and
(4.24) $\lambda\geq\frac{n+1}{2}\kappa.$
Moreover, if the identity in (4.24) holds, then $(M,\omega_{g})$ is isometrically biholomorphic to a ball quotient with constant holomorphic bisectional curvature.
The following well-known maximum principle (e.g. [HL11, Theorem 2.10]) is used in the proof of Theorem 4.4.
Lemma 4.3.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded domain, and $L$ be a uniformly elliptic second order differential operator with continuous coefficients. If $u\in C^{2}(\Omega)\cap C^{0}(\overline{\Omega})$ satisfies $Lu\geq 0$ and $u\leq 0$, then either $u<0$ or $u\equiv 0$.
Theorem 4.4.
Let $(M,\omega_{g})$ be a compact Hermitian manifold of dimension $n$, $(N,\omega_{h})$ be a Kähler manifold, and $f:M\to N$ be a non-constant holomorphic map. If there exist two constants $\lambda>0$ and $\kappa>0$ such that
(4.25) $\mathrm{Ric}^{\mathrm{(2)}}(\omega_{g})\geq\lambda f^{*}\omega_{h}\quad\mbox{and}\quad R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)\leq\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h},$
for any $v\in T^{1,0}M$, then
(4.26) $\mathrm{tr}_{\omega_{g}}f^{*}\omega_{h}\leq\frac{n(n+1)\kappa}{2\lambda}.$
Moreover, if the identity in (4.26) holds at some point $p\in M$, then $f^{*}\omega_{h}=\frac{(n+1)\kappa}{2\lambda}\omega_{g}$ and $(M,\omega_{g})$ is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},2\kappa^{-1}\omega_{\mathrm{FS}}\right)$.
Proof.
By using similar arguments as in the proof of Theorem 1.14, we have
(4.27) $\Box_{g}u\geq\frac{\lambda}{n}u^{2}-\frac{(n+1)\kappa}{2}u$
where $u=\mbox{tr}_{\omega_{g}}f^{*}\omega_{h}$. Since $M$ is compact, we may assume $M$ attains its maximum at $p\in M$. Therefore
$0\geq\Box_{g}u(p)\geq\frac{\lambda}{n}u(p)^{2}-\frac{(n+1)\kappa}{2}u(p),$
and we get
$\sup_{M}u=u(p)\leq\frac{n(n+1)\kappa}{2\lambda}.$
This gives (4.26). Suppose the identity in (4.26) holds at some point $p\in M$. We set
$\widetilde{u}=u-\frac{n(n+1)\kappa}{2\lambda}.$
Therefore, $\widetilde{u}\leq 0$ and $\widetilde{u}(p)=0$. A straightforward computation shows that
(4.28) $\Box_{g}\widetilde{u}=\Box_{g}u\geq\frac{\lambda}{n}u\left(u-\frac{n(n+1)\kappa}{2\lambda}\right)\geq\frac{(n+1)\kappa}{2}\widetilde{u}.$
By using the maximum principle in Lemma 4.3, we deduce that $\widetilde{u}\equiv 0$. Therefore,
(4.29) $u\equiv\frac{n(n+1)\kappa}{2\lambda}.$
In this case, one can see clearly that $\nabla df=0$,
(4.30) $\left(g^{i\overline{j}}R^{g}_{i\overline{j}k\overline{\ell}}\right)g^{k\overline{q}}g^{p\overline{\ell}}h_{\alpha\overline{\beta}}f^{\alpha}_{p}\overline{f^{\beta}_{q}}=\frac{\lambda}{n}u^{2},\ \ \ R^{h}_{\alpha\overline{\beta}\gamma\overline{\delta}}\left(g^{i\overline{j}}f^{\alpha}_{i}\overline{f^{\beta}_{j}}\right)\left(g^{p\overline{q}}f^{\gamma}_{p}\overline{f^{\delta}_{q}}\right)=\frac{(n+1)\kappa}{2}u.$
This implies
$\mathrm{Ric}^{\mathrm{(2)}}(\omega_{g})=\lambda f^{*}\omega_{h}\quad\mbox{and}\quad R^{h}\left(f_{*}v,\overline{f_{*}v},f_{*}v,\overline{f_{*}v}\right)=\kappa|v|^{2}_{g}|f_{*}v|^{2}_{h}.$
By using similar arguments as in the proof Theorem 1.16, one has
$f^{*}\omega_{h}=\frac{(n+1)\kappa}{2\lambda}\omega_{g}.$
Hence, $g$ is a Kähler metric with constant holomorphic sectional curvature $\kappa$, and so $(M,\omega_{g})$ is isometrically biholomorphic to $\left({\mathbb{C}}{\mathbb{P}}^{n},2\kappa^{-1}\mathrm{\omega_{\mathrm{FS}}}\right)$. ∎
5. Rigidity of holomorphic maps between vector bundles
In this section, we prove Theorem 1.17 and Theorem 1.20.
Proof of Theorem 1.17. We are supposing by contradiction that $\varphi$ is non-trivial. By using the same setup as in the proof of Theorem 2.1, we get the estimate (2.11) at point $x_{0}\in M$, i.e., for any $w\in T^{1,0}_{x_{0}}M$,
(5.1) $0\geq\left(\partial\overline{\partial}\log F\right)(w,\overline{w})\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
Since $(E_{1},G)$ is $k$-RC positive, there exists a $k$-dimensional linear subspace $W$ of $T_{x_{0}}^{1,0}M$ such that for all $w\in W\setminus\{0\}$
$R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)>0.$
If $W\cap\ker df\neq\{0\}$, then there exists a non-zero $w\in W$ such that $f_{*}w=0$. Plugging this $w$ into (5.1), one obtains
(5.2) $0\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}>0,$
and this is a contradiction. Hence, $W\cap\ker df=\{0\}$, and $df(W)$ is a $k$-dimensional linear subspace of $T_{f(x_{0})}^{1,0}N$. Since $(E_{2},H)$ is $s$-RC non-positive and $k+s>\dim N$, there exists some non-zero $w\in W$ such that
$R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)>0\quad\mbox{and}\quad R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)\leq 0.$
Plugging this $w$ into (5.1), one obtains
(5.3) $0\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}>0,$
and this is again a contradiction. Therefore, $\varphi$ must be a trivial map. ∎
Proof of Theorem 1.20. Suppose by contradiction that $\varphi$ is non-trivial. We use the same setup as in the proof of Theorem 2.1. In particular, we have the estimate (2.11) at point $x_{0}\in M$, i.e., for any $w\in T^{1,0}_{x_{0}}M$,
(5.4) $0\geq\left(\partial\overline{\partial}\log F\right)(w,\overline{w})\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{G}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{H}}.$
Since $(E_{1},G)$ is $k$-RC non-negative, there exists a $k$-dimensional linear subspace $W$ of $T_{x_{0}}^{1,0}M$ such that for all $w\in W$
$R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)\geq 0.$
Let $r_{0}$ be the complex rank of $df$ at $x_{0}$. One can see clearly that
$\dim df(W)\geq\dim W-\dim\ker\left(df(x_{0})\right)=k-(\dim M-r_{0})\geq k+r-\dim M.$
Since $(E_{2},H)$ is $s$-RC negative and $\left(k+r-\dim M\right)+s>\dim N$, there exists some non-zero $w\in W$ such that
$R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)\geq 0\quad\mbox{and}\quad R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)<0.$
Plugging this $w$ into (5.4), one obtains
(5.5) $0\geq\frac{R^{E_{1}}\left(w,\overline{w},\xi,\overline{\xi}\right)}{|\xi|_{g}^{2}}-\frac{R^{E_{2}}\left(f_{*}w,\overline{f_{*}w},\varphi(\xi),\overline{\varphi(\xi)}\right)}{|\varphi(\xi)|^{2}_{h}}>0,$
and this is a contradiction. Therefore, $\varphi$ must be a trivial map. ∎
By using similar arguments as in the proof of Theorem 1.17, one can prove the following results (see e.g. [Yang21], [Yang24]).
Corollary 5.1.
Let $(M,\omega_{g})$ be a compact Hermitian manifold and $(N,\omega_{h})$ be a Hermitian manifold. If
  1. (1)
    $(M,\omega_{g})$ is RC-positive; and
  2. (2)
    $(N,\omega_{h})$ has non-positive holomorphic bisectional curvature,
then any holomorphic map $f:M\rightarrow N$ is a constant map.
Corollary 5.2.
Let $(M,\omega_{g})$ be a compact Hermitian manifold and $(N,\omega_{h})$ be a Hermitian manifold. If
  1. (1)
    $(M,\omega_{g})$ has $\mathrm{HSC}>0$ (resp. $\mathrm{HSC}\geq 0$); and
  2. (2)
    $(N,\omega_{h})$ has $\mathrm{HSC}\leq 0$ (resp. $\mathrm{HSC}<0$),
then any holomorphic map $f:M\rightarrow N$ is a constant map.